# Random Sperner lemma and random Brouwer fixed point theorem

### v3

@@ 2.1 Different forms and special cases of random Brouwer fixed point theorem
$\bar{L}^{0}(\mathcal{F})$ denotes the set of equivalence classes of extended real-valued random variables on $(\Omega,\mathcal{F},P)$. For any $\xi,\eta\in\bar{L}^{0}(\mathcal{F})$, we define $\xi\leq\eta$ iff $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ for almost all $\omega$ in $\Omega$ (briefly, $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ a.s.), where $\xi^{0}$ and $\eta^{0}$ are arbitrarily chosen representatives of $\xi$ and $\eta$, respectively, then $(\bar{L}^{0}(\mathcal{F}),\leq)$ forms a complete lattice (see [12] for details).
As usual, $\xi<\eta$ means $\xi\leq\eta$ and $\xi\neq\eta$, whereas, for any $A\in\mathcal{F}$, $\xi<\eta$ on $A$ means $\xi^{0}(\omega)<\eta^{0}(\omega)$ for almost all $\omega\in A$, where $\xi^{0}$ and $\eta^{0}$ are again arbitrary representatives of $\xi$ and $\eta$, respectively. Furthermore, we denote $L_{+}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi\geq 0\}$ and $L_{++}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi>0~{}on~{}\Omega\}$. For any $A\in\mathcal{F}$, we use $\tilde{I}_{A}$ for the equivalence class of $I_{A}$, where $I_{A}$ is the characteristic function of $A$ (namely $I_{A}(\omega)=1$ if $\omega\in A$ and $I_{A}(\omega)=0$ otherwise).
@@
[…]
@@
  • $\sigma$-stable if $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}\in G$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ (namely each $A_{n}\in\mathcal{F},A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$, and $\bigcup_{n=1}^{\infty}A_{n}=\Omega$). Here, $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$ is well defined, namely it has a natural explanation: arbitrarily choosing a representative $x_{n}^{0}$ of $x_{n}$ for each $n\in\mathbb{N}$, the function $\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0}$ defined by $(\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0})(\omega)=x_{n}^{0}(\omega)$ when $\omega\in A_{n}$ for some $n\in\mathbb{N}$, is exactly a representative of $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$.
@@
  • a.s. sequentially continuous if for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ such that $\{\|x_{n}-x\|,n\in\mathbb{N}\}$ converges a.s. to 00, it holds that $\{\|f(x_{n})-f(x)\|,n\in\mathbb{N}\}$ converges a.s. to 00.
  • random sequentially continuous if $G$ is $\sigma$-stable and for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ convergent in $\mathcal{T}_{\varepsilon,\lambda}$ to $x$ there exists a random subsequence $\{x_{n_{k}},k\in\mathbb{N}\}$ of $\{x_{n},n\in\mathbb{N}\}$ such that $\{f(x_{n_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(x)$. Here, $\{x_{n_{k}},k\in\mathbb{N}\}$ is a random subsequence of $\{x_{n},n\in\mathbb{N}\}$, we means $\{n_{k},k\in\mathbb{N}\}$ is a sequence of equivalence classes of positive-integer-valued random variables such that the following two conditions are satisfied:
@@
One of the main purposes of this paper is to provide a complete proof of the following random Brouwer fixed point theorem — Theorem 2.1, which was given a sketch proof in [26] (see Lemma 4.6 of [26]).
Theorem 2.1.
@@
It is obvious that a $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping defined on a $\sigma$-stable set is random sequentially continuous, Proposition 2.3 below, which first occurred in [28, Proposition 4.2] is a special case of Proposition 2.2.
@@
As usual, $2^{\mathbb{R}^{d}}$ denotes the family of subsets of $\mathbb{R}^{d}$. Let $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ be a set-valued mapping satisfying the following two conditions: (1) $U(\omega)$ is a nonempty bounded closed convex subset of $\mathbb{R}^{d}$ for each $\omega\in\Omega$; (2) $Gr(U):=\{(\omega,x)\in\Omega\times\mathbb{R}^{d}:x\in U(\omega)\}\in\mathcal{F}\otimes Bor(\mathbb{R}^{d})$, namely $U$ has a measurable graph, where $Bor(\mathbb{R}^{d})$ stands for the Borel $\sigma$-algebra of $\mathbb{R}^{d}$.
Further, for $U$ as above, we always denote by $G_{U}$ the set of equivalence classes of almost surely measurable selections of $U$, namely $G_{U}:=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)~{}a.s.\}$.
Then it is not very difficult to verify that $G_{U}$ is an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (see Lemma 5.3 of this paper), so $G_{U}$ is also $\sigma$-stable and it is easy to check that a mapping $h:G_{U}\rightarrow G_{U}$ is local iff $h$ is $\sigma$-stable (see Lemma 5.1 of this paper).
Proposition 2.4 below, which was called the stochastic Brouwer fixed point theorem in [44] and has played an essential role in the proofs of Theorems 5.1 and 5.2 of [45]. It is obvious that Proposition 2.4 is a special case of Proposition 2.3. The original proof of Proposition 2.4 was based on Ponosov’s solution of the Nemytskii conjecture [42] and a measurable selection theorem from [51]. In this paper, we can even prove that Proposition 2.4 is equivalent to Proposition 2.3 by means of Ponosov’s result [42]. In particular, we can also obtain a somewhat surprising conclusion again by means of Ponosov’s result [42]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous, see Proposition 5.8 of this paper.
Proposition 2.4 is a slightly more general formulation of Theorem 2.1 of [44] (see also Lemma B.4 of [45]) in that we remove the assumption on completeness of $(\Omega,\mathcal{F},P)$ (otherwise, we first consider $(\Omega,\hat{\mathcal{F}},P)$, which is the completion of $(\Omega,\mathcal{F},P)$, and obtain a fixed point $x$ of $h$ as in [44] and $x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)$ a.s. and $x^{0}$ is $\hat{\mathcal{F}}$-measurable, whereas $x^{0}$ is equal a.s. to an $\mathcal{F}$-measurable function $\hat{x}^{0}$, then the equivalence class $\hat{x}$ of $\hat{x}^{0}$ is exactly our desired fixed point).
@@
A very special case of Proposition 2.3 first appeared as Proposition 3.1 in [11], which states that every $\sigma$-stable a.s. sequentially continuous mapping from an a.s. bounded a.s. sequentially closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point, where a subset is said to be a.s. sequentially closed if it contains all a.s. limits of sequences in it. Proposition 2.3 is a generalization of Proposition 3.1 of [11] since a.s. sequentially closed set is exactly $\mathcal{T}_{\varepsilon,\lambda}$-closed and a.s. sequential continuity obviously implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. It is interesting to notice that Proposition 5.8 implies the equivalence of Proposition 3.1 of [11] with either of Proposition 2.3 and Proposition 2.4.
We would also like to point out that Proposition 2.2 (namely Theorem 2.1) and Proposition 2.3 are both best possible, which is explained as follows. It is well known that $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a metrizable linear topological space under $\mathcal{T}_{\varepsilon,\lambda}$ (namely the topology of convergence in probability), there is, naturally, the notion of boundedness in the sense of the linear topology $\mathcal{T}_{\varepsilon,\lambda}$, denoted by the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, it is easy to see that the a.s. boundedness implies the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, and we have known from Proposition 16 of [24] that these two kinds of boundedness coincide for a $\sigma$-stable subset. Likewise, the $L^{0}$-convexity implies the usual convexity, and we have also known from Lemma 2.8 of [49] that these two kinds of convexity coincide for a $\mathcal{T}_{\varepsilon,\lambda}$-closed and $\sigma$-stable subset. On the one hand, an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is generally noncompact, if the hypothesis on the $\sigma$-stability of $f$ in Proposition 2.3 is removed, then Proposition 2.3 does not hold, which comes from a classical result due to Dobrowolski and Marciszewski [10] (see also [26, Section 2.1]). On the other hand, an example provided by Ponosov in [44] shows that Proposition 2.4 (namely Proposition 2.3) does not hold if $G_{U}$ is replaced with a generic $\mathcal{T}_{\varepsilon,\lambda}$-closed, $\mathcal{T}_{\varepsilon,\lambda}$-bounded and convex set $G$ (namely $G$ is not $L^{0}$-convex). Therefore, they are all best possible.
@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
In this paper, we present a general theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes and give a thorough treatment of random Sperner lemma by grasping the inherent connection between $L^{0}$-simplexes and classical simplexes, so we can smoothly present a good definition of an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (see Definition 2.6 below) and establish a general random Sperner lemma (see Theorem 2.9 below). To survey these main results, let us first introduce some necessary notation and prerequisites.
@@
For any $A,B\in\mathcal{F}$, we will always use the corresponding lowercase letters $a$ and $b$ for the equivalence classes $[A]$ and $[B]$ (two elements $C$ and $D$ in $\mathcal{F}$ are said to be equivalent if $P[(C\setminus D)\cup(D\setminus C)]=0$), respectively. Define $1=[\Omega]$, $0=[\emptyset]$, $a\wedge b=[A\cap B]$, $a\vee b=[A\cup B]$ and $a^{c}=[A^{c}]$, where $A^{c}$ denotes the complement of $A$. Let $B_{\mathcal{F}}=\{a=[A]:A\in\mathcal{F}\}$, then $(B_{\mathcal{F}},\wedge,\vee,^{c},0,1)$ is a complete Boolean algebra, namely a complete complemented distributive lattice (see [34] for details). Specially, $B_{\mathcal{F}}$ is called the measure algebra associated with $(\Omega,\mathcal{F},P)$.
As usual, for any $a,b\in B_{\mathcal{F}}$, $a>b$ means $a\geq b$ and $a\neq b$. A subset $\{a_{i}:i\in I\}$ of $B_{\mathcal{F}}$ is called a partition of unity if $\bigvee_{i\in I}a_{i}=1$ and $a_{i}\wedge a_{j}=0$ for any $i,j\in I$ with $i\neq j$. The collection of all such partitions is denoted by $p(1)$. For any $\{a_{j}:j\in J\}\in p(1)$, it is clear that the set $\{j\in J:a_{j}>0\}$ is at most countable. Moreover, $\{a_{n},n\in\mathbb{N}\}\in p(1)$ iff there exists a countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ such that $a_{n}=[A_{n}]$ for each $n\in\mathbb{N}$.
@@
Since our future target is to develop random functional analysis and its applications to dynamic Nash equilibrium theory and nonsmooth differential geometry on metric measure spaces (see [4, 5, 6, 16, 17, 36, 37] for recent advances), random normed modules and random locally convex modules will play an active role and they are all more general $L^{0}$-modules than $L^{0}(\mathcal{F},\mathbb{R}^{d})$, and thus we state the results of this section in the $L^{0}$-modules for possible convenience in the future applications.
In the special $L^{0}$-module $L^{0}(\mathcal{F},\mathbb{R}^{d})$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ is always well defined for any sequence $\{x_{n},n\in\mathbb{N}\}$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$. However, for a generic $L^{0}$-module $E$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ may be meaningless for any sequence $\{x_{n},n\in\mathbb{N}\}$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $E$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$, but once a nonempty subset $G$ of $E$ is $\sigma$-stable, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ is well defined for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$. The notion of a $\sigma$-stable set was introduced in [23] (see also (4) of Definition 2.5 below) in order to make sense the countable concatenation such as $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$.
@@
    A finite subset $\{x_{1},x_{2},\cdots,x_{n}\}$ of $E$ is said to be $L^{0}$-affinely independent if either $n=1$ or if $n>1$ and $\{x_{i}-x_{1}:i=2\sim n\}$ is $L^{0}$-linearly independent, in which case $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ is called an $(n-1)$-dimensional $L^{0}$-simplex with vertexes $x_{1},x_{2},\cdots,x_{n}$ and $Conv_{L^{0}}(\{x_{i}:i\in I\})$ is called an $(|I|-1)$-dimensional $L^{0}$-face of $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ for any nonempty subset $I$ of $\{1,2,\cdots,n\}$, where $|I|$ denotes the cardinality of $I$.

### v5

@@ 2.1 The random Brouwer fixed point theorem and its special cases
$\bar{L}^{0}(\mathcal{F})$ denotes the set of equivalence classes of extended real-valued random variables on $(\Omega,\mathcal{F},P)$. For any $\xi,\eta\in\bar{L}^{0}(\mathcal{F})$, a partial order $\leq$ on $\bar{L}^{0}(\mathcal{F})$ is defined as follows: $\xi\leq\eta$ iff $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ for almost all $\omega$ in $\Omega$ (briefly, $\xi^{0}(\omega)\leq\eta^{0}(\omega)$ a.s.), where $\xi^{0}$ and $\eta^{0}$ are arbitrarily chosen representatives of $\xi$ and $\eta$, respectively. Then $(\bar{L}^{0}(\mathcal{F}),\leq)$ forms a complete lattice (see [13] for details).
As usual, $\xi<\eta$ means $\xi\leq\eta$ and $\xi\neq\eta$, whereas, for any $A\in\mathcal{F}$, $\xi<\eta$ on $A$ means $\xi^{0}(\omega)<\eta^{0}(\omega)$ for almost all $\omega\in A$, where $\xi^{0}$ and $\eta^{0}$ are again arbitrary representatives of $\xi$ and $\eta$, respectively. Furthermore, we denote $L_{+}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi\geq 0\}$ and $L_{++}^{0}(\mathcal{F}):=\{\xi\in L^{0}(\mathcal{F}):\xi>0~on~\Omega\}$. For any $A\in\mathcal{F}$, we use $\tilde{I}_{A}$ for the equivalence class of $I_{A}$, where $I_{A}$ is the characteristic function of $A$ (namely, $I_{A}(\omega)=1$ if $\omega\in A$ and $I_{A}(\omega)=0$ otherwise).
@@
For any $x\in L^{0}(\mathcal{F},\mathbb{R}^{d})$ and any $r\in L_{++}^{0}(\mathcal{F})$, the set $B(x,r):=\{y\in L^{0}(\mathcal{F},\mathbb{R}^{d}):\|x-y\|<r~\text{on}~\Omega\}$ is called the random open ball centered at $x$ with the random radius $r$, where the $L^{0}$-norm $\|\cdot\|$ is defined by $\|x\|=\sqrt{\sum_{i=1}^{d}x_{i}^{2}}$ for any $x=(x_{1},x_{2},\cdots,x_{d})\in L^{0}(\mathcal{F},\mathbb{R}^{d})$. The collection $\{B(x,r):x\in L^{0}(\mathcal{F},\mathbb{R}^{d}),r\in L_{++}^{0}(\mathcal{F})\}$ forms a base for some Hausdorff topology on $L^{0}(\mathcal{F},\mathbb{R}^{d})$. Consequently, $L^{0}(\mathcal{F},\mathbb{R}^{d})$ can be equipped with two different topologies: the topology of convergence in probability, denoted by $\mathcal{T}_{\varepsilon,\lambda}$, and the topology generated by random open balls, denoted by $\mathcal{T}_{c}$. $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is called the random Euclidean space of rank $d$ when it is endowed with either $\mathcal{T}_{\varepsilon,\lambda}$ or $\mathcal{T}_{c}$.
@@
  • $\sigma$-stable if $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}\in G$ for any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ and any countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ (namely, each $A_{n}\in\mathcal{F},A_{i}\cap A_{j}=\emptyset$ for any $i\neq j$, and $\bigcup_{n=1}^{\infty}A_{n}=\Omega$). Here, $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$ is well defined, namely, it allows a natural interpretation: arbitrarily choosing a representative $x_{n}^{0}$ of $x_{n}$ for each $n\in\mathbb{N}$, the function $\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0}$ defined by $(\sum_{n=1}^{\infty}I_{A_{n}}x_{n}^{0})(\omega)=x_{n}^{0}(\omega)$ when $\omega\in A_{n}$ for some $n\in\mathbb{N}$, is exactly a representative of $\sum_{n=1}^{\infty}\tilde{I}_{A_{n}}x_{n}$.
@@
  • a.s. sequentially continuous if for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ such that $\{\|x_{n}-x\|,n\in\mathbb{N}\}$ converges a.s. to 0, then we have $\{\|f(x_{n})-f(x)\|,n\in\mathbb{N}\}$ converges a.s. to 0.
  • random sequentially continuous if $G$ is $\sigma$-stable and for any $x\in G$ and any sequence $\{x_{n},n\in\mathbb{N}\}$ in $G$ convergent in $\mathcal{T}_{\varepsilon,\lambda}$ to $x$ there exists a random subsequence $\{x_{n_{k}},k\in\mathbb{N}\}$ of $\{x_{n},n\in\mathbb{N}\}$ such that $\{f(x_{n_{k}}),k\in\mathbb{N}\}$ converges in $\mathcal{T}_{\varepsilon,\lambda}$ to $f(x)$. Here, by $\{x_{n_{k}},k\in\mathbb{N}\}$ being a random subsequence of $\{x_{n},n\in\mathbb{N}\}$, we mean that $\{n_{k},k\in\mathbb{N}\}$ is a sequence of equivalence classes of positive-integer-valued random variables such that the following two conditions are satisfied:
@@
Theorem 2.1 below is just the so called the random Brouwer fixed point theorem (Earlier, a sketch of proof, rather incomplete, was given in [27, Lemma 4.6]).
Theorem 2.1 (The Random Brouwer Fixed Point Theorem).
@@
It is obvious that a $\mathcal{T}_{\varepsilon,\lambda}$-continuous mapping defined on a $\sigma$-stable set is random sequentially continuous. Proposition 2.3 below, which was first given in [29, Proposition 4.2] is a special case of Proposition 2.2 (namely, Theorem 2.1).
@@
As usual, $2^{\mathbb{R}^{d}}$ denotes the family of subsets of $\mathbb{R}^{d}$. Let $U:\Omega\rightarrow 2^{\mathbb{R}^{d}}$ be a set-valued mapping satisfying the following two conditions: (1) $U(\omega)$ is a nonempty bounded closed convex subset of $\mathbb{R}^{d}$ for each $\omega\in\Omega$; (2) $Gr(U):=\{(\omega,x)\in\Omega\times\mathbb{R}^{d}:x\in U(\omega)\}\in\mathcal{F}\otimes Bor(\mathbb{R}^{d})$, namely, $U$ has a measurable graph, where $Bor(\mathbb{R}^{d})$ stands for the Borel $\sigma$-algebra of $\mathbb{R}^{d}$.
Further, for $U$ as above, we always denote by $G_{U}$ the set of equivalence classes of almost surely measurable selections of $U$, namely, $G_{U}:=\{x\in L^{0}(\mathcal{F},\mathbb{R}^{d}):x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)~a.s.\}$.
Then it is not difficult to verify that $G_{U}$ is an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ (see Lemma 5.3 of this paper), so $G_{U}$ is also $\sigma$-stable and it is easy to check that a mapping $h:G_{U}\rightarrow G_{U}$ is local iff $h$ is $\sigma$-stable (see Lemma 5.1 of this paper).
Proposition 2.4 below, which was called the stochastic version of Brouwer fixed point theorem in [46], has played an essential role in the proofs of Theorems 5.1 and 5.2 of [47]. It is obvious that Proposition 2.4 is a special case of Proposition 2.3. The original proof of Proposition 2.4 was based on Ponosov’s solution of the Nemytskii conjecture [44] and a measurable selection theorem from [55]. In this paper, we can even prove that Proposition 2.4 is equivalent to Proposition 2.3 by means of Ponosov’s result [44]. In particular, we can also obtain a somewhat surprising conclusion again by means of Ponosov’s result [44]: a $\mathcal{T}_{\varepsilon,\lambda}$-continuous $\sigma$-stable mapping from a $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to $L^{0}(\mathcal{F},\mathbb{R}^{d})$ must be a.s. sequentially continuous, see Proposition 5.8 of this paper.
Proposition 2.4 is a slightly more general formulation of Theorem 2.1 of [46] (see also Lemma B.4 of [47]) in that we have removed the assumption on completeness of $(\Omega,\mathcal{F},P)$ (otherwise, we first consider $(\Omega,\hat{\mathcal{F}},P)$, which is the completion of $(\Omega,\mathcal{F},P)$, and obtain a fixed point $x$ of $h$ as in [46] and $x$ has a representative $x^{0}$ such that $x^{0}(\omega)\in U(\omega)$ a.s. and $x^{0}$ is $^\hat{\mathcal{F}}$-measurable, whereas $x^{0}$ is equal a.s. to an $\mathcal{F}$-measurable function $\hat{x}^{0}$. Then the equivalence class $\hat{x}$ of $\hat{x}^{0}$ is exactly our desired fixed point).
@@
A very special case of Proposition 2.3 first appeared as Proposition 3.1 in [12], which states that every $\sigma$-stable a.s. sequentially continuous mapping from an a.s. bounded a.s. sequentially closed $L^{0}$-convex subset of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ to itself has a fixed point, where a subset is said to be a.s. sequentially closed if it contains all a.s. limits of sequences in it. Proposition 2.3 is a generalization of Proposition 3.1 of [12] since a.s. sequentially closed set is exactly $\mathcal{T}_{\varepsilon,\lambda}$-closed and a.s. sequential continuity obviously implies $\mathcal{T}_{\varepsilon,\lambda}$-continuity. It is interesting to notice that Proposition 5.8 implies the equivalence of Proposition 3.1 of [12] with either of Proposition 2.3 or Proposition 2.4.
We would also like to point out that Proposition 2.2 (namely, Theorem 2.1) and Proposition 2.3 are both best possible, which is explained as follows. It is well known that $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is a metrizable linear topological space under $\mathcal{T}_{\varepsilon,\lambda}$ (namely, the topology of convergence in probability). There is, naturally, the notion of boundedness in the sense of the linear topology $\mathcal{T}_{\varepsilon,\lambda}$, denoted by the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness. It is easy to see that the a.s. boundedness implies the $\mathcal{T}_{\varepsilon,\lambda}$-boundedness, and we have known from Proposition 16 of [25] that these two kinds of boundedness coincide for a $\sigma$-stable subset. Likewise, the $L^{0}$-convexity implies the usual convexity, and we have also known from Lemma 2.8 of [52] that these two kinds of convexity coincide for a $\mathcal{T}_{\varepsilon,\lambda}$-closed and $\sigma$-stable subset. On the one hand, an a.s. bounded $\mathcal{T}_{\varepsilon,\lambda}$-closed $L^{0}$-convex subset $G$ of $L^{0}(\mathcal{F},\mathbb{R}^{d})$ is generally noncompact. If the hypothesis on the $\sigma$-stability of $f$ in Proposition 2.3 is removed, then Proposition 2.3 does not hold, which comes from a classical result due to Dobrowolski and Marciszewski [11] (see also [27, Section 2.1]). On the other hand, an example provided by Ponosov in [46] shows that Proposition 2.4 (namely, Proposition 2.3) does not hold if $G_{U}$ is replaced with a generic $\mathcal{T}_{\varepsilon,\lambda}$-closed, $\mathcal{T}_{\varepsilon,\lambda}$-bounded and convex set $G$ (namely, $G$ is not $L^{0}$-convex). Therefore, they are all best possible.
@@ 2.2 A general $L^{0}$-simplicial subdivision and random Sperner lemma
In this paper, we present a general theory of $L^{0}$-simplicial subdivisions of $L^{0}$-simplexes and give a thorough treatment of a random Sperner lemma by utilizing the inherent connection between $L^{0}$-simplexes and classical simplexes. So we can smoothly present a good definition of an $L^{0}$-simplicial subdivision of an $L^{0}$-simplex (see Definition 2.6 below) and establish a general random Sperner lemma (see Theorem 2.9 below). To survey these main results, we first introduce some notation and prerequisites.
@@
For any $A,B\in\mathcal{F}$, we will always use the corresponding lowercase letters $a$ and $b$ for the equivalence classes $[A]$ and $[B]$ (two elements $C$ and $D$ in $\mathcal{F}$ are said to be equivalent if $P[(C\setminus D)\cup(D\setminus C)]=0$), respectively. Define $1=[\Omega]$, $0=[\emptyset]$, $a\wedge b=[A\cap B]$, $a\vee b=[A\cup B]$ and $a^{c}=[A^{c}]$, where $A^{c}$ denotes the complement of $A$. Let $B_{\mathcal{F}}=\{a=[A]:A\in\mathcal{F}\}$, then $(B_{\mathcal{F}},\wedge,\vee,^{c},0,1)$ is a complete Boolean algebra, namely, a complete complemented distributive lattice (see [35] for details). Specifically, $B_{\mathcal{F}}$ is called the measure algebra associated with $(\Omega,\mathcal{F},P)$.
As usual, for any $\in B_{\mathcal{F}}, a>ba>b$ means $a\geq b$ and $a\neq b$. A subset $\{a_{i}:i\in I\}$ of $B_{\mathcal{F}}$ is called a partition of unity if $\bigvee_{i\in I}a_{i}=1$ and $a_{i}\wedge a_{j}=0$ for any $i,j\in I$ with $i\neq j$. The collection of all such partitions is denoted by $p(1)$. For any $\{a_{j}:j\in J\}\in p(1)$, it is clear that the set $\{j\in J:a_{j}>0\}$ is at most countable. Moreover, $\{a_{n},n\in\mathbb{N}\}\in p(1)$ iff there exists a countable partition $\{A_{n},n\in\mathbb{N}\}$ of $\Omega$ to $\mathcal{F}$ such that $a_{n}=[A_{n}]$ for each $n\in\mathbb{N}$.
@@
Since our aim is to develop random functional analysis and its applications to dynamic Nash equilibrium theory and nonsmooth differential geometry on metric measure spaces (see [4, 6, 5, 17, 18, 37, 38] for recent advances), random normed modules will play an active role and they are all more general $L^{0}$-modules than $L^{0}(\mathcal{F},\mathbb{R}^{d})$, and thus we state the results of this section in terms of $L^{0}$-modules for possible convenience in future applications.
In the special $L^{0}$-module $L^{0}(\mathcal{F},\mathbb{R}^{d})$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ is always well defined for any sequence $\{x_{n},n\in\mathbb{N}\}$ and any $\{a_{n},n\in\mathbb{N}\}\in p(1)$. However, for a generic $L^{0}$-module $E$, $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ may be meaningless for an arbitrary sequence $\{x_{n},n\in\mathbb{N}\}$ in $E$ and arbitrary $\{a_{n},n\in\mathbb{N}\}\in p(1)$. For this, we need the notion of a $\sigma$-stable set introduced in [24] (see also (4) of Definition 2.5 below) in order to speak of a countable concatenation such as $\sum_{n=1}^{\infty}I_{a_{n}}x_{n}$ in a meaningful way.
@@
    A finite subset $\{x_{1},x_{2},\cdots,x_{n}\}$ of $E$ is said to be $L^{0}$-affinely independent if either $n=1$ or if $n>1$ and $\{x_{i}-x_{1}:i=2\sim n\}$ is $L^{0}$-linearly independent, in which case $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ is called an (n1)(n-1)-dimensional $L^{0}$-simplex with vertices $x_{1},x_{2},\cdots,x_{n}$ and $Conv_{L^{0}}(\{x_{i}:i\in I\})$ is called an (|I|1)(|I|-1)-dimensional $L^{0}$-face of $Conv_{L^{0}}(\{x_{1},x_{2},\cdots,x_{n}\})$ for any nonempty subset $I$ of $\{1,2,\cdots,n\}$, where $|I|$ denotes the cardinality of $I$.